# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v3

@@ E.1 Common Setup for All FairPO Variants
Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison. The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k. During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features. All experiments were conducted on the MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009] datasets. Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops. The AdamW optimizer [Loshchilov and Hutter, 2019] was used to update all trainable parameters. Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set. The reference model parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, required for FairPO-DPO and the non-privileged loss constraint in all variants, were obtained from a BCE-SFT model detailed in Appendix D.
E.2 Per-Label Non-Linear MLP Classifier Head
@@
E.3 FairPO Variant-Specific Hyperparameters
The core FairPO learning rates, $\eta_{\mathbf{w}}$ for model parameters and $\eta_{\alpha}$ for GRPO weights, were tuned via grid search. For $\eta_{\mathbf{w}}$ (the AdamW LR), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$, and for $\eta_{\alpha}$, values were explored from $\{0.01,0.05,0.1,0.2\}$. After tuning, the final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$. The specific hyperparameters for each FairPO variant’s privileged loss component (e.g., $\beta$, $\gamma$, $\lambda_{\text{CPO}}$) were also tuned on the designated validation set for each dataset, with final values reported in Table 5 (Appendix F). Throughout all experiments, the GRPO mechanism with scaled loss updates, as described in Section 3, was used to adaptively balance the privileged and non-privileged loss terms.
Appendix F Hyperparameter Details
F.1 Hyperparameters Tuning
For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019] with a batch size of 32. The base model was a vit-base-patch16-224222https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k. The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$, while the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$. All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP, using a patience of 5 epochs [^3] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
Specific hyperparameters for each FairPO variant were tuned via grid search on the validation set. For FairPO-DPO, the strength parameter $\beta$ was chosen from $\{0.1,0.5,1.0\}$, and the non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$. For FairPO-SimPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the margin $\gamma$ from $\{0.05,0.1,0.2\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. For FairPO-CPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the NLL regularizer weight $\lambda_{\text{CPO}}$ from $\{0.1,0.5,1.0\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
For our baselines, the loss weight for privileged labels in BCE-SFT + Privileged Re-Weighting was set to 5 after tuning from $\{2,3,5,10\}$. The group weight learning rate $\eta_{\alpha}$ for GDRO + BCE was tuned similarly to FairPO.
F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE. Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$. Base learning rate $\eta_{\mathbf{w}}$ was 1e41e-4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
Dataset Method $\beta$ / $\lambda_{\text{CPO}}$ $\gamma$ (SimPO only) $\epsilon$
MS-COCO FairPO-DPO $\beta=0.5$ N/A $0.05$
FairPO-SimPO $\beta=0.5$ $0.1$ $0.05$
FairPO-CPO $\beta=0.5,\lambda_{\text{CPO}}=0.5$ N/A $0.01$
NUS-WIDE FairPO-DPO $\beta=0.1$ N/A $0.05$
FairPO-SimPO $\beta=0.1$ $0.05$ $0.1$
FairPO-CPO $\beta=0.1,\lambda_{\text{CPO}}=0.5$ N/A $0.05$
Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values. Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results.
Table 6: Sensitivity of FairPO-CPO mAP on MS-COCO to preference strength $\beta$ (with $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$).
$\beta$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.05 88.95 90.28
0.1 89.32 90.31
0.5 (Optimal) 89.76 90.34
1.0 89.51 90.25
2.0 88.67 90.11
Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed. A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
Table 7: Sensitivity of FairPO-CPO mAP on MS-COCO to non-privileged slack $\epsilon$ (with $\beta=0.5,\lambda_{\text{CPO}}=0.5$).
$\epsilon$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.001 89.65 90.40
0.01 (Optimal) 89.76 90.34
0.05 89.81 90.22
0.1 89.70 90.05
0.2 89.55 89.87

### v4

@@ G.1 Common Setup for All FairPO Variants
Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison. The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 (dosovitskiy2021imageworth16x16words), which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k. During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features. All experiments were conducted on the MS-COCO 2014 (lin2015microsoftcococommonobjects) and NUS-WIDE (chua2009nus) datasets. Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops. The AdamW optimizer (loshchilov2019decoupled) was used to update all trainable parameters. Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set.
G.2 Per-Label Non-Linear MLP Classifier Head